OpenStudy (anonymous):

Find the indefinite integral integral (3t^2)/(t^2+1) dt

hartnn (hartnn):

did you try substituting t = tan x ?

OpenStudy (anonymous):

I got 3(tan^(-1)(t))+C

hartnn (hartnn):

but i got 3 (t - tan^(-1)t) + c can you show your work, so that i couls spot the error ?

hartnn (hartnn):

numerator = tan^2 x = 1+ sec^2 x because of that "sec^2 x", i got its integral as tan x, which is just t

hartnn (hartnn):

sorry, tan^2x = sec^2 x -1 :P

OpenStudy (mathmale):

Another approach you might want to consider: Note that the numerator and denom. of the given integrand are of the same order (same power of t). DIVIDE the den. into the num. BEFORE attempting to integrate.

OpenStudy (anonymous):

ok thanks.S

hartnn (hartnn):

welcome ^_^